OpenStudy (rayford):

(1). Solve 3^x – 4 = 7^x + 9 x ≈ –25.86 x ≈ –2.59 x ≈ –0.39 x ≈ –0.04 (2). Solve 5^x + 5 = 9^x x ≈ –3.63 x ≈ –0.28 x ≈ 0.07 x ≈ 13.69

OpenStudy (anonymous):

those choices are approximations so i'm assuming that you need to solve by graphing (calculator)?

OpenStudy (anonymous):

no. not really if you look at the problem you will notice there are a like terms so this should break down and you don't need a graphing calc for this problem

OpenStudy (anonymous):

to further show what I am talking because we have 3^x -4 +=7x^x +9 in this case we can see that the power of x is one because no value is given.

OpenStudy (anonymous):

the answer should be given like the answers shown

OpenStudy (anonymous):

ingore the extra x on 7^X it is a typo

OpenStudy (jiteshmeghwal9):

\(\LARGE3^x-9=7^x+4\)

OpenStudy (jiteshmeghwal9):

\(\LARGE3^x-3^3=7^x+4\)

mathslover (mathslover):

@jiteshmeghwal9 one mistake it will be 3^2

OpenStudy (jiteshmeghwal9):

\(\LARGE3^{x \over 2}=7^x+9\) Ohhhh ! sorry thanx @mathslover :)

OpenStudy (jiteshmeghwal9):

4*

mathslover (mathslover):

No problem continue jitesh

OpenStudy (jiteshmeghwal9):

\(\LARGE3^{x \over 2}=7^x+4\)

OpenStudy (jiteshmeghwal9):

here comes the right problem :(

OpenStudy (jiteshmeghwal9):

wht should we do further @mathslover @waterineyes ????

OpenStudy (jiteshmeghwal9):

need ur help :)

OpenStudy (anonymous):

\[3^x - 7^x = 13\]

mathslover (mathslover):

\[\large{ 3^x – 4 = 7^x + 9}\] \[\large{3^x-7^x=9+4}\]

OpenStudy (jiteshmeghwal9):

so wht further ??? Need to solve for 'x'

mathslover (mathslover):

http://www.wolframalpha.com/input/?i=solve+%3A3%5Ex-7%5Ex%3D9%2B4 wolfram has no step by step soln also

OpenStudy (jiteshmeghwal9):

i also find out this http://www.wolframalpha.com/input/?i=3^x-4%3D7^x%2B9

OpenStudy (anonymous):

We have to go by plugging in the values from the answer choices and using calculator that is obvious..

mathslover (mathslover):

we can not do by plugging in values since it don't have real soln

OpenStudy (jiteshmeghwal9):

well if we turn bases same then can we do it as \(3^x=9\) \(3^x=3^2\) \(x=2\) ?????

mathslover (mathslover):

No no , i think there is involvement of logarithm here

OpenStudy (jiteshmeghwal9):

hmmmm

OpenStudy (anonymous):

We cannot evaluate log for : \(log(a + b)\)..

OpenStudy (anonymous):

Is there any formula for this ??

mathslover (mathslover):

log(a + b) = log(a * (1 + b/a)) = log a + log(1 + b/a)

mathslover (mathslover):

log(3^x) + log(1+7^x/3^x)

mathslover (mathslover):

\[\large{xlog3 +log(\frac{3^x+7^x}{3^x})}\]

mathslover (mathslover):

\[\large{x*log3 + log(3^+7^x)-log(3^x)}\] \[\large{\cancel{xlog3}+log(3^x+7^x)\cancel{-log(3^x)}}\] Proved :D

mathslover (mathslover):

@vishweshshrimali5

OpenStudy (anonymous):

How will you find : \[\log(1 + \frac{b}{a})\]

OpenStudy (anonymous):

They are still in addition..

mathslover (mathslover):

simple : \[\large{log(\frac{a+b}{a})}\] \[\large{log(a+b)-log(a)}\]

OpenStudy (anonymous):

And here how will you find: \[\log(a + b)\]

mathslover (mathslover):

that is the main problem

OpenStudy (anonymous):

They are again addition..

OpenStudy (anonymous):

Ha ha ha.. That is what I am saying..

mathslover (mathslover):

:)

OpenStudy (vishweshshrimali5):

@waterineyes @mathslover and everyone else... well I am preparing for a competitive exam which has MCQs so I have got a sort of expertise in them and one of the way of solving questions which will be difficult to solve (like thisone) we usually plug in the values of the options and find out the correct answer and simply evaluate them using our common sense/.......... I think this will be the best method(for I have not tried solving it till now) for solving this question

OpenStudy (anonymous):

I was saying the same..

mathslover (mathslover):

but @vishweshshrimali5 the given values are too complex

mathslover (mathslover):

would you like to find : 3^(-25.6) or anything else?

mathslover (mathslover):

Or we are going to do that by guess?

OpenStudy (vishweshshrimali5):

we can use binomial theorem for 3^(x/2) and 7^(x)

OpenStudy (anonymous):

Try smaller ones.. Ha ha ha..

mathslover (mathslover):

.............

OpenStudy (vishweshshrimali5):

we can write \(7^x = (3+4)^x\)

mathslover (mathslover):

:D @waterineyes

OpenStudy (vishweshshrimali5):

and then simplify them.......

mathslover (mathslover):

:'(

OpenStudy (vishweshshrimali5):

for answer no. 1 its obvious that for 3^x - 7^x = 13 any real value of x is not going to satisfy it.............

mathslover (mathslover):

right

OpenStudy (vishweshshrimali5):

and for 2nd 9^x - 5^x = 5 now its clear that for x<0 ;9^x <5^x so that would not work

OpenStudy (vishweshshrimali5):

so only 2 options remain

OpenStudy (vishweshshrimali5):

for x = 0.07 its clear that both 9^x and 5^x will be very very small to get their difference equal to 5

OpenStudy (vishweshshrimali5):

only option that remain is the last one

OpenStudy (vishweshshrimali5):

@waterineyes

OpenStudy (vishweshshrimali5):

The both questions are WRONG

OpenStudy (vishweshshrimali5):

Because 9 - 5 = 4 which is very close to 5 thus its clear that x should be close 1

OpenStudy (vishweshshrimali5):

So, the correct answer should be 1

OpenStudy (vishweshshrimali5):

@rayford you just fooled all of us ........ XD

OpenStudy (vishweshshrimali5):

@waterineyes got my point ?????????

OpenStudy (vishweshshrimali5):

@jiteshmeghwal9

OpenStudy (anonymous):

Yeah I got @vishweshshrimali5

OpenStudy (vishweshshrimali5):

thanks for the medal everyone :) gud @waterineyes

OpenStudy (rayford):

i'm sorry i fell asleep last night... so what answers would you suggest choosing vishweshshrimali5?

OpenStudy (rayford):

or someone else..

OpenStudy (vishweshshrimali5):

@rayford hi See, as I have already mentioned the both questions are wrong, so there is no chance of saying which option is correct

OpenStudy (vishweshshrimali5):

for question 1, there is no such real value of x. for question 2, the value of x to the nearest integer = 1

OpenStudy (rayford):

just throwing this out there..but you do realize the problem looks like this right?

OpenStudy (rayford):